OpenStudy (sphott51):

3(x-5)^1/3 = -9 the answer is {-22} needs to show the work, and dont know how to get there;

OpenStudy (freckles):

first isolate the (x-5)^(1/3) by undoing the multiplication there by 3 to undo that multiplication by 3 you will have to divide both sides by 3 that is do: \[\frac{3(x-5)^\frac{1}{3}}{3}=\frac{-9}{3} \\ \frac{3}{3}(x-5)^\frac{1}{3}=\frac{-9}{3}\] I will let you finish this step

OpenStudy (freckles):

are you there?

OpenStudy (anonymous):

step by step:

OpenStudy (anonymous):

3(x-5)^1/3 = -9 (divide both sides by 3) (-9/3=-3) (x-5)^1/3 = -3 (raise both sides to exponent 3 to ger riid of cube root 1/3) [(x-5)^1/3]^3 = (-3)^3 (exponent 3 cancels exponent 1/3 since 3/3=1) (x-5) = -27 (take -5 to other side) x = -22